# Toxicity Detection is NOT all you Need: Measuring the Gaps to Supporting Volunteer Content Moderators through a User-Centric Method

### v2

@@ Appendix C User Survey
No Ambiguous/ Misleading/ Inaccurate Information or Clickbait in The Submission Title - misinformation
misinformation
TwHIN-BERT-Misinformation-Classifier
No
No
Yes, but with changes
No
Detect misinformation for Twitter posts - 1. Need to be adapted to title-like texts. 2. Need to be adapted to a movie-related context
No Ambiguous/ Misleading/ Inaccurate Information or Clickbait in The Submission Title - Clickbait
clickbait
distilroberta-clickbait
No
No
Yes
Yes
No Extraneous Comic Book Movie Submission
misinformation
TwHIN-BERT-Misinformation-Classifier
No
No
Yes, but with changes
No
Detect misinformation for Twitter posts - Need to be adapted to a movie-related context
No Repost
rule-based
Yes
No
No Discussion Threads of New Releases
rule-based
Yes
No
No Spoilers
spoiler
roberta-base-finetuned-imdb-spoilers
No
No
Yes
Yes
No Encouraging Piracy
piracy
Classification
No
No
Yes, but with changes
No
Detect if a website is a piracy website - Need to detect not only piracy website but also comments that encourage piracy
No Subreddit Brigading
brigading
NA - Toxicity
No
Yes
Yes, but with changes
No
These models may be able to detect some of the brigading comments, but there is no specific model for brigading and some brigading encouraging comments may not be caught by the toxicity models
Prohibited Popular Websites
rule-based
Yes
No
@@
Civility
toxicity; hatespeech
Text-Moderation
No
Yes
Yes
The model may not be able to catch Holocaust denialism, which is vital in this case.
Current event
time/year prediction for event; history event; event extractor
GoLLIE-7B
No
No
Yes, but with changes
Information extraction
1. The model needs to be adapted to question-contexts and extract event names from the question. 2. Then by matching the event name with Wikipedia page, we may get the year or time of the event.
Homework
homework
NA
No
No
No
Scope
topic classifier
tweet-topic-21-multi
No
No
Yes, but with changes
Twitter topic classification
Need to be fine-tuned to this topic.
NSFW/Sensitive content
NSFW; sexual
NSFW_text _classifier
No
No
Yes
"Poll-type" question
poll
NA
No
No
No
"Soapboxing" or loaded questions
soapboxing; agenda pushing
NA
No
No
No
No "what if" questions
hypothetical; counterfactual; alternative history
NA
No
No
No
Privacy
PII; personal private
lgpd_pii_ identifier
No
No
Yes, but with changes
Identify sensitive data in the scope of LGPD. The goal is to have a tool to identify document numbers like CNPJ, CPF, people’s names, and other kinds of sensitive data, allowing companies to find and anonymize data according to their business needs, and governance rules.
Need to be adapted to the sensitive data here
Illegal artifacts
illegal; safety
beaver-dam-7b
No
No
Yes
Basic facts
basic facts; factoid
nf-cats
No
No
Yes
Depth
depth
NA
No
No
No
Sources
reliable source
NA
No
No
No
No personal anecdotes
personal anecdote
muppet-roberta-base-joke_detector
No
No
Yes, but with changes
Detect jokes, stories, and anecdotes
The model is trained with 2000 jokes. Here the detection target is not about jokes. In addition, here the need is focusing on personal anecdotes.
No speculation
speculation; factual; fact-check
verdict-classifier-en
No
No
Yes, but with changes
Fact-checking model for detecting misinformation
The model does not detect speculation or personal opinion. Here the detection target is the speculation of history events, which can be especially hard for such a fact-checking model.
No partial answers or "placeholders"
partial answer
NA
No
No
No
No political agendas or moralizing
political agenda; moralizing
NA
No
No
No
Do not just post links or quotations
rule-based
NA
Yes
No
No plagiarism
plagiarism
longformer-base-plagiarism-detection
No
No
Yes, but with changes
Detect machine-paraphrased plagiarisms
The real-world source of plagiarism is way larger than the training and testing dataset size of this model.
Jokes and humor
joke; humor; flippant
muppet-roberta-base-joke_detector
No
No
Yes
Digression
relevance
response-quality-classifier-large
No
No
Yes, but with changes
Rate the relevancy and specificity of a response within a dialogue
Need to be adapted to the Reddit (question-comment) texts.
Clutter
clutter
NA
No
No
No

### v3

@@ Appendix D User Survey
No Ambiguous/ Misleading/ Inaccurate Information or Clickbait in The Submission Title - misinformation misinformation TwHIN-BERT-Misinformation-Classifier No No Yes, but with changes No Detect misinformation for Twitter posts - 1. Need to be adapted to title-like texts. 2. Need to be adapted to a movie-related context
No Ambiguous/ Misleading/ Inaccurate Information or Clickbait in The Submission Title - Clickbait clickbait distilroberta-clickbait No No Yes Yes
No Extraneous Comic Book Movie Submission misinformation TwHIN-BERT-Misinformation-Classifier No No Yes, but with changes No Detect misinformation for Twitter posts - Need to be adapted to a movie-related context
No Repost rule-based Yes No
No Discussion Threads of New Releases rule-based Yes No
No Spoilers spoiler roberta-base-finetuned-imdb-spoilers No No Yes Yes
No Encouraging Piracy piracy Classification No No Yes, but with changes No Detect if a website is a piracy website - Need to detect not only piracy website but also comments that encourage piracy
No Subreddit Brigading brigading NA - Toxicity No Yes Yes, but with changes No These models may be able to detect some of the brigading comments, but there is no specific model for brigading and some brigading encouraging comments may not be caught by the toxicity models
Prohibited Popular Websites rule-based Yes No
@@
Civility toxicity; hatespeech Text-Moderation No Yes Yes The model may not be able to catch Holocaust denialism, which is vital in this case.
Current event time/year prediction for event; history event; event extractor GoLLIE-7B No No Yes, but with changes Information extraction 1. The model needs to be adapted to question-contexts and extract event names from the question. 2. Then by matching the event name with Wikipedia page, we may get the year or time of the event.
Homework homework NA No No No
Scope topic classifier tweet-topic-21-multi No No Yes, but with changes Twitter topic classification Need to be fine-tuned to this topic.
NSFW/Sensitive content NSFW; sexual NSFW_text _classifier No No Yes
"Poll-type" question poll NA No No No
"Soapboxing" or loaded questions soapboxing; agenda pushing NA No No No
No "what if" questions hypothetical; counterfactual; alternative history NA No No No
Privacy PII; personal private lgpd_pii_ identifier No No Yes, but with changes Identify sensitive data in the scope of LGPD. The goal is to have a tool to identify document numbers like CNPJ, CPF, people’s names, and other kinds of sensitive data, allowing companies to find and anonymize data according to their business needs, and governance rules. Need to be adapted to the sensitive data here
Illegal artifacts illegal; safety beaver-dam-7b No No Yes
Basic facts basic facts; factoid nf-cats No No Yes
Depth depth NA No No No
Sources reliable source NA No No No
No personal anecdotes personal anecdote muppet-roberta-base-joke_detector No No Yes, but with changes Detect jokes, stories, and anecdotes The model is trained with 2000 jokes. Here the detection target is not about jokes. In addition, here the need is focusing on personal anecdotes.
No speculation speculation; factual; fact-check verdict-classifier-en No No Yes, but with changes Fact-checking model for detecting misinformation The model does not detect speculation or personal opinion. Here the detection target is the speculation of history events, which can be especially hard for such a fact-checking model.
No partial answers or "placeholders" partial answer NA No No No
No political agendas or moralizing political agenda; moralizing NA No No No
Do not just post links or quotations rule-based NA Yes No
No plagiarism plagiarism longformer-base-plagiarism-detection No No Yes, but with changes Detect machine-paraphrased plagiarisms The real-world source of plagiarism is way larger than the training and testing dataset size of this model.
Jokes and humor joke; humor; flippant muppet-roberta-base-joke_detector No No Yes
Digression relevance response-quality-classifier-large No No Yes, but with changes Rate the relevancy and specificity of a response within a dialogue Need to be adapted to the Reddit (question-comment) texts.
Clutter clutter NA No No No